# Dynamics and fluid-structure interaction in turbulent flows within and above flexible canopies

### 21 Dec 2023 v1

@@ 3.1 Mean flow quantities
To investigate the flow above and within the canopy for different values of $Ca$, we start looking at the mean profiles of the velocity and of the Reynolds stresses, along with relevant derived quantities. In the following, averaging in time and along the homogeneous directions is denoted with angle brackets, while fluctuations with respect to such mean are marked with an apostrophe. The mean profiles of the streamwise velocity $u$, shown in panel $a$ of figure 3, exhibit two distinct inflection points: an outer inflection point (square symbol) generated by the drag discontinuity at the average canopy tip position and an inner inflection point (triangle symbol) closer to the wall, where the inflected velocity profile connects to the wall boundary layer. Below the outer inflection point the inner flow is reminiscent of that attained in an anisotropic porous medium (Rosti et al., 2018b), while above that the mean velocity profile of the outer flow is similar to that of a turbulent boundary layer over a rough wall (Monti et al., 2019). We therefore compute the virtual origin of the outer flow, $y_{vo}$, imposing the matching with a canonical logarithmic profile and notice that, for all the cases of interest, it lays well between the two inflection points (panel $b$ of figure 3), thus confirming that we are in a dense canopy regime (Monti et al., 2020, 2022). The canopy becomes more compliant to the flow increasing $Ca$, hence the mean position of its tip as well as all the other relevant points are monotonously shifted downward, maintaining their relative order; the mean streamwise velocity at those points, instead, exhibits a non monotonous trend. At the canopy tip, in particular, it first undergoes a slight increase due to the reduction of the filaments drag, later decreasing again as the effect of the downward shift becomes dominant. The positions of all points reach a plateau for high values of $Ca$, where the vertical stacking of the deflected filaments poses a lower bound to the thickness of the inner flow region. As expected, on scaling the velocity profile of the inner flow with the friction velocity computed at the wall, $u_{\tau}^{in}=\sqrt{\tau_{w}/\rho_{f}}=\sqrt{\nu\partial\langle u\rangle/\partial y|_{0}}$, the typical trend ${\langle u\rangle}/{u_{\tau}^{in}}=yu_{\tau}^{in}/\nu$ is recovered close to the bottom wall (on the left of panel $a$, figure 4). Instead, on scaling the velocity profile of the outer flow with the friction velocity computed at the virtual origin, $u_{\tau}^{out}=\sqrt{\nu\partial\langle u\rangle/\partial y|_{y_{vo}}-\langle u^{\prime}v^{\prime}\rangle|_{y_{vo}}}$, we confirm good agreement with a logarithmic profile (on the right of panel $a$, figure 4) of the form
@@
where $\kappa=0.41$ and $B=5.2$, while $\Delta u^{+}_{out}$ denotes the friction function accounting for the mean velocity shift in the outer flow due to the presence of the canopy (or wall roughness, as in Jiménez, 2004). As noted by Monti et al. (2022), $\Delta u^{+}_{out}$ exhibits an exponential trend with the driving pressure gradient $\mathrm{d}P/\mathrm{d}x$ (panel $b$ of figure 4), once made dimensionless upon the height of the channel above the virtual origin. We acknowledge a minor deviation from their data, that we impute to the different shape of our filaments and to their movement.
Figure 3: Mean profiles of the streamwise velocity (panel $a$) for different values of $Ca$ and associated relevant points (inset and panel $b$). In panel $b$ we show the position of the relevant points for different values of $Ca$, maintaining the the vertical scale unchanged with respect to that of panel $a$.
Figure 4: Panel $a$ reports the mean velocity profiles made dimensionless upon the friction velocity at the wall ($u_{\tau}^{*}=u_{\tau}^{in}$) / at the virtual origin ($u_{\tau}^{*}=u_{\tau}^{out}$) against the wall distance shifted by $y_{w}^{*}=0$ / $y_{w}^{*}=y_{vo}$. The inner scaling yields good overlapping of the profiles with a linear trend over the first grid points off the wall while, with the outer scaling, the profiles collapse on those computed with equation 5, denoted with a dash-dotted line. The friction function $\Delta u^{+}_{out}$ appearing in equation 5 has an exponential trend with respect to the driving pressure gradient, $\mathrm{d}P/\mathrm{d}x$, as reported in panel $b$ and confirmed by the data of Monti et al. (2022).
@@ 3.2 Energy spectra
In panels $a$ and $b$ of figure 8 we inspect the profiles of $E_{x}$ and $E_{z}$, respectively, for different values of $Ca$ at the inner inflection point, $y=y_{in}$, at the outer inflection point, $y=y_{tip}$, and above the canopy, at $y=H/2$. We also report the typical $-5/3$ slope for comparison. Within the canopy, both spectra exhibit a non-monotonous behaviour characterised by a first peak at the wave number associated to the channel height and a second, sharper one, associated to the mean separation of the filaments. In between the two, the spectra approach a $-5/3$ slope when $Ca$ increases. The filaments absorb energy in the shear production range, close to the largest scales of motion, and re-inject it in the flow through their wakes and their waving motion, in close agreement to the spectral short cut process described by Finnigan (2000) and Olivieri et al. (2020). A variation in the $Ca$ is observed to affect both the spectral short cut and the amplitude of the spectra. In facts, the spectral short cut is more intense in the case of rigid filaments, where the spectra almost approach a plateau before their second peak, while it appears less accentuated in the cases at higher $Ca$, consistently with the observations of Olivieri et al. (2022a). A more regular decay of the spectra is recovered moving up to the drag discontinuity at the canopy tip. In particular, while $E_{x}$ decreases monotonously from the largest scales of motion, $E_{z}$ peaks at $k_{z}=\mathcal{O}(10)$: this behaviour, which is also found in the outer flow at $y=H/2$, appears compatible with the presence of large structures dominating the outer flow above the canopy (Monti et al., 2023). Conventional full spectra are observed in the outer flow, denoting the persistence of a fully-developed turbulent state there. Interestingly, the amplitude of the spectra within the canopy appears to increase with $Ca$, while an opposite trend is observed outside. An increase in the flexibility of the filaments is in facts associated to more intense velocity fluctuations at the inner inflection point, located deep in the canopy and marginally affected from the outer flow turbulence, due to the increased motion of the filaments. At and above the canopy tip, instead, more flexible filaments are associated to the generation of a weaker shear layer with respect to the rigid case due to the lower drag discontinuity at their tip.
@@ 3.3 Lumley triangle
Our analysis supports a multi-layer approach to the modelling of canopy flow turbulence like in Poggi et al. (2004), based on the collection of different turbulence states along the wall normal direction. Turbulence in the inner flow exhibits a peculiar structure close to the bottom wall, tending towards a more isotropic condition immediately above and further transitioning smoothly towards a state similar to that attained in the log-law region of turbulent wall flows. The profiles of $\xi$ and $\eta$ show a reasonable collapse in both the most rigid ($Ca=0,1,10$) and most flexible cases ($Ca=50,100$), once scaled in inner viscous units, with the case at $Ca=25$ laying in between the two. Turbulence immediately above the canopy is more isotropic than over the no-slip wall of the open channel ($\eta$ is smaller), consistently with the observations of Kuwata & Suga (2016) for an isotropic porous medium. Instead, the portion of the outer flow from the log-law region up to where the upper free-slip wall is felt appears analogous to a turbulent wall flow, thus suggesting outer similarity arguments. There, the anisotropy settles on an almost linear trend common to all $Ca$.
@@ 3.4 Inner / outer flow interaction
a)b)
@@
Figure 10: Instantaneous sweep and ejection events at the tip of a rigid (panel $a$) and a flexible (panel $b$, $Ca=100$) canopy. The regions where the events are occurring are delimited with black lines, while their magnitude is quantified as $|uw|/U_{b}^{2}$ and visualised with a linear colormap ranging from white to orange (ejections) or violet (sweeps) in [0,0.4].
@@
The previous subsection was devoted to the characterisation of the inner and outer flows, which we separately investigated in order to emphasise their individual features. Here, instead, we focus on the interaction between the flow within and above the canopy, thus turning our attention to the events taking place close to the canopy tip. As extensively discussed in literature, the flows over plant canopies (Finnigan, 2000) along with those above permeable (Breugem & Boersma, 2005) and elastic (Rosti & Brandt, 2017) walls are often dominated by large scale vortical structures elongated in the spanwise direction (rollers), generated by a Kelvin-Helmholtz like instability. The rollers induce large scale sweep ($u^{\prime}>0,v^{\prime}<0$) and ejection ($u^{\prime}<0,v^{\prime}>0$) events, visualised in figure 10, promoting momentum transfer at the canopy-fluid interface. In order to provide a quantitative description of those events, we compute the joint probability density function (J-PDF) of the streamwise and wall-normal velocity fluctuations and sample it at the average position of the canopy tip for different values of $Ca$. As visible in figure 11, the peaks always lay in the second quadrant ($u^{\prime}<0,v^{\prime}>0$), thus denoting the dominance of ejections over sweeps. The downward motion of high-speed fluid is mitigated by the presence of the canopy, which instead opposes little resistance to the uplift of low-speed fluid from its interior. Such observation is consistent with the experiment of Chowdhuri et al. (2022), who noted that long-lasting ejection events are more relevant than short-lived sweep events close to the canopy tip. Notwithstanding the similarities, the J-PDFs exhibit different shapes due to the effect of canopy flexibility. In the most rigid cases the J-PDFs elongate in the fourth quadrant ($u^{\prime}>0,v^{\prime}<0$), remarking the occurrence of rare and intense sweep events which are attenuated increasing the flexibility. In the most flexible cases, in facts, the filaments are deflected by the mean flow and shield the inner region of the canopy from downward interactions. Furthermore, at $Ca=0$ and $Ca=1$, the J-PDFs exhibit a secondary peak in the third quadrant ($u^{\prime}<0,v^{\prime}<0$) associated with the deflection of the streamlines around the filament tip. As seen in Monti et al. (2020) and in figure 16 of Nicholas et al. (2023), immediately after the tip the flow decelerates and plunges.
@@
Our analysis appears to support a scenario where turbulence inside the canopy is sustained by the spanwise rollers in the outer flow. Those are accompanied by intense sweeps and ejections, but while the former are hindered by the presence of the filaments, the canopy opposes little resistance to the latter. The filaments, in facts, regulate the vertical exchange of streamwise momentum allowing for the transpiration of the canopy in correspondence of the low pressure regions accompanying the spanwise rollers and shielding the inner flow from all but the most intense downward interactions, associated to the secondary instability of the rollers (Nepf, 2012a). Such intense sweeps induce a significant local deflection of the filaments and a large-amplitude flapping motion. Turbulence inside the canopy is therefore likely driven by frequent ejections and rare intense sweep events (Nepf, 2012a), along with the flapping motion of the filaments in the flexible cases.

### 14 May 2024 v2

@@ 3.1 Mean flow quantities
To investigate the flow above and within the canopy for different values of $Ca$, we start looking at the mean profiles of the velocity and of the Reynolds stresses, along with relevant derived quantities. In the following, averaging in time and along the homogeneous directions is denoted with angle brackets, while fluctuations with respect to such mean are marked with an apostrophe. Our analysis is based on $100$ flow fields for each value of $Ca$, regularly collected over $50$ bulk time units. The flow statistics do not show any appreciable variation upon computation with half of the fields, thus confirming that they are converged.
@@
The mean profiles of the streamwise velocity $u$, shown in panel $a$ of figure 3, exhibit two distinct inflection points: an outer inflection point ($y_{out}$, square symbol) generated by the drag discontinuity at the average canopy tip position and an inner inflection point ($y_{in}$, triangle symbol) closer to the wall, where the inflected velocity profile connects to the wall boundary layer. Below the outer inflection point the inner flow is reminiscent of that attained in an anisotropic porous medium (Rosti et al., 2018b), while immediately above that the outer flow is similar to a turbulent mixing layer (Raupach et al., 1996). Yet, canopy flows can also be considered instances of obstructing substrates for the assessment of outer-layer similarity (Chen & García-Mayoral, 2023). Fitting a logarithmic profile to the mean velocity well above the canopy is in facts a convenient numerical expedient to simplify its parametrisation for modelling purposes, even though it does not satisfactorily represent the physical features of the flow. We therefore compute the virtual origin of the outer flow, $y_{vo}$, imposing the matching with a canonical logarithmic profile and notice that, for all the cases of interest, it lays well between the two inflection points (panel $b$ of figure 3), thus confirming that we are in a dense canopy regime (Monti et al., 2020, 2022). The canopy becomes more compliant to the flow increasing $Ca$, hence the mean position of its tip as well as all the other relevant points are monotonously shifted downward, maintaining their relative order; the mean streamwise velocity at those points, instead, exhibits a non monotonous trend. At the canopy tip, in particular, it first undergoes a slight increase due to the reduction of the filaments drag, later decreasing again as the effect of the downward shift becomes dominant. The positions of all points reach a plateau for high values of $Ca$, where the vertical stacking of the deflected filaments poses a lower bound to the thickness of the inner flow region.
As expected, on scaling the velocity profile of the inner flow with the friction velocity computed at the wall, $u_{\tau}^{in}=\sqrt{\tau_{w}/\rho_{f}}=\sqrt{\nu\partial\langle u\rangle/\partial y|_{0}}$, the typical trend ${\langle u\rangle}/{u_{\tau}^{in}}=yu_{\tau}^{in}/\nu$ is recovered close to the bottom wall (panel $a$ of figure 4). Instead, on scaling the velocity profile of the outer flow with the friction velocity computed at the virtual origin, $u_{\tau}^{out}=\sqrt{\nu\partial\langle u\rangle/\partial y|_{y_{vo}}-\langle u^{\prime}v^{\prime}\rangle|_{y_{vo}}}$, we confirm good agreement with a logarithmic profile (panel $b$ of figure 4) of the form
@@
where $\kappa=0.41$ and $B=5.2$, while $\Delta u^{+}_{out}$ denotes the friction function accounting for the mean velocity shift in the outer flow due to the presence of the canopy (or wall roughness, as in Jiménez, 2004). As noted by Monti et al. (2022), to whom we compare our results, $\Delta u^{+}_{out}$ exhibits an exponential trend with the driving pressure gradient $\mathrm{d}P/\mathrm{d}x$ (panel $c$ of figure 4), once made dimensionless upon the height of the channel above the virtual origin. We acknowledge a minor deviation from their data, that we impute to the different shape of our filaments and to their movement. Differently from our, in facts, the study of Monti et al. (2022) only concerned rigid canopies: in particular, at a chosen value of the Reynolds number ($Re_{b}=6000$, higher than ours), they varied the canopy solidity $\lambda$ by changing the inclination of the filaments.
Figure 3: Mean profiles of the streamwise velocity (panel $a$) for different values of $Ca$ and associated relevant points (inset and panel $b$). In panel $b$ we show the position of the relevant points for different values of $Ca$, maintaining the vertical scale unchanged with respect to that of panel $a$.
Figure 4: Panel $a$ reports the first five computed points of the mean velocity profiles, made dimensionless upon the friction velocity at the wall, $u_{\tau}^{in}$, against the wall distance. Panel $b$, instead, reports the mean velocity profiles made dimensionless upon the friction velocity at the virtual origin, $u_{\tau}^{out}$, against the wall distance shifted by $y_{vo}$, within the region where the scaling holds. In panel $b$, for each case, we also report the logarithmic profile computed with equation 5 as a dash-dotted gray line. The inner scaling yields good overlapping of the different profiles with a quasi-linear trend over the first grid points off the wall while, with the outer scaling, the profiles collapse on the analytical predictions. Finally, in panel $c$, we show the exponential trend of the friction function $\Delta u^{+}_{out}$ appearing in equation 5 with respect to the driving pressure gradient, $\mathrm{d}P/\mathrm{d}x$, and compare it with the numerical data of Monti et al. (2022), who studied rigid canopies with different inclinations.
@@
3.2 Canopy drag
Here we focus on the measurement of the canopy drag coefficient, $C_{d}$. First, in close analogy to Ghisalberti & Nepf (2006), we define
$C_{d}a(y)=\displaystyle\frac{\displaystyle\frac{\partial\langle u^{\prime}v^{\prime}\rangle}{\partial y}\bigg{\rvert}_{y_{out}<y<H}-\frac{\partial\langle u^{\prime}v^{\prime}\rangle}{\partial y}(y)}{\displaystyle\frac{1}{2}\langle u\rangle^{2}(y)}$ (7)
where $a$ is the frontal canopy area per unit volume, and the first term in the numerator denotes the mean vertical gradient of the Reynolds’ shear stress above the canopy tip, $y_{out}$. We thus compute the mean value of $C_{d}a$, denoted with angle brackets, between the inner and the outer inflection points for different values of $Ca$. This approach circumvents the complexity of the computation of $a$ (Ghisalberti & Nepf, 2006), but yields a dimensional quantity which we thus make dimensionless upon multiplication with the channel height. The outcome is compared with the experimental measurements of Ghisalberti & Nepf (2006) in panel $a$ of figure 8, highlighting a good correspondence between the two for the most flexible cases considered in our investigation. The data appear to tend towards $\langle C_{d}a\rangle H\sim Ca^{-1/4}$ at large values of $Ca$.
Figure 8: Canopy drag measurements. We show (panel $a$) the mean value of $C_{d}a$ from our simulations, compared to that measured experimentally by Ghisalberti & Nepf (2006). We also report (panel $b$) the value of the canopy drag, $C_{d}$, throughout our simulations.
In order to decouple the value of $C_{d}$ from $a$, we consider the integral form of the streamwise momentum balance
[…]
where the streamwise pressure gradient $\frac{dP}{dx}$ equates the sum of the mean shear stress at the wall, $\tau_{w}$, and of the canopy drag, $D_{c}(y)$, integrated across the canopy height. $C_{d}$ can thus be computed through such balance, referring it to the frontal canopy area,
[…]
The outcome of this approach in our simulations is reported in panel $b$ of figure 8. The value of $C_{d}$ in the rigid case is compatible with what reported in literature for canopies with similar properties (Raupach & Thom, 1981; Shimizu et al., 1992; Finnigan, 2000; Ghisalberti & Nepf, 2006; Nepf, 2012a), where $C_{d}\approx 0.5-1.5$, and quickly decreases reducing the rigidity of the filaments, reaching a plateau for $Ca\gtrapprox 50$. Overall, the drag measurements reported here are in good agreement with the experimental measurements and shed light on the trend of $C_{d}$ with $Ca$.
@@ 3.3 Energy spectra
In panels $a$ and $b$ of figure 9 we inspect the profiles of $E_{x}$ and $E_{z}$, respectively, for different values of $Ca$ at the inner inflection point, $y=y_{in}$, at the outer inflection point, $y=y_{tip}$, and above the canopy, at $y=H/2$. We also report the typical $-5/3$ slope for comparison. Within the canopy, both spectra exhibit a non-monotonous behaviour characterised by a first peak at the wave number associated to the channel height and a second, sharper one, associated to the mean separation of the filaments. In between the two, the spectra approach a $-5/3$ slope when $Ca$ increases. The filaments absorb energy in the shear production range, close to the largest scales of motion, and re-inject it in the flow through their wakes and their waving motion, in close agreement to the spectral short cut process described by Finnigan (2000) and Olivieri et al. (2020). A variation in the $Ca$ is observed to affect both the spectral short cut and the amplitude of the spectra. In facts, the spectral short cut is more intense in the case of rigid filaments, where the spectra almost approach a plateau before their second peak, while it appears less accentuated in the cases at higher $Ca$, consistently with the observations of Olivieri et al. (2022a). A more regular decay of the spectra is recovered moving up to the drag discontinuity at the canopy tip. In particular, while $E_{x}$ decreases monotonously from the largest scales of motion, $E_{z}$ peaks at $k_{z}\sim\mathcal{O}(10)$: this behaviour, which is also found in the outer flow at $y=H/2$, appears compatible with the presence of large structures dominating the outer flow above the canopy (Monti et al., 2023). Conventional full spectra are observed in the outer flow, denoting the persistence of a fully-developed turbulent state there. Interestingly, the amplitude of the spectra within the canopy appears to increase with $Ca$, while an opposite trend is observed outside. An increase in the flexibility of the filaments is in facts associated to more intense velocity fluctuations at the inner inflection point, located deep in the canopy and marginally affected from the outer flow turbulence, due to the increased motion of the filaments. At and above the canopy tip, instead, more flexible filaments are associated to the generation of a weaker shear layer with respect to the rigid case due to the lower drag discontinuity at their tip.
@@ 3.4 Lumley triangle
Our analysis supports a multi-layer approach to the modelling of canopy flow turbulence like in Poggi et al. (2004), based on the collection of different turbulence states along the wall normal direction. Turbulence in the inner flow exhibits a peculiar structure close to the bottom wall, tending towards a more isotropic condition immediately above and further transitioning smoothly towards a state similar to that attained in the log-law region of turbulent wall flows. This picture is also supported by the visualisations of the instantaneous, local anisotropy of the flow reported in appendix A. The profiles of $\xi$ and $\eta$ in figure 10 show a reasonable collapse in both the most rigid ($Ca=0,1,10$) and most flexible cases ($Ca=50,100$), once scaled in inner viscous units, with the case at $Ca=25$ laying in between the two. Turbulence immediately above the canopy is more isotropic than over the no-slip wall of the open channel ($\eta$ is smaller), consistently with the observations of Kuwata & Suga (2016) for an isotropic porous medium. Instead, the portion of the outer flow from the log-law region up to where the upper free-slip wall is felt appears analogous to a turbulent wall flow, thus suggesting outer similarity arguments. There, the anisotropy settles on an almost linear trend common to all $Ca$.
@@ 3.5 Inner / outer flow interaction
a)Refer to captionb)Refer to captionc)
d)Refer to captione)Refer to captionf)
@@
Figure 11: Instantaneous sweep and ejection events in a rigid (panels $a,b,c$) and a flexible (panels $d,e,f$) canopy flow with $Ca=100$, at $y=H/2$ (panels $a,d$), at the canopy tip (panels $b,e$) and at the virtual origin (panels $c,f$). The flow is sampled on wall-parallel planes with the mean velocity aligned to the vertical direction, going from bottom to top. Regions where the events are occurring are delimited with black lines, while their magnitude is quantified as $|uw|/U_{b}^{2}$ and visualised with a linear colormap ranging from white to orange (ejections) or violet (sweeps) in [0,0.4].
Multiple authors have proposed layered descriptions of canopy flows, identifying regions with specific dynamical properties while moving along the wall-normal direction. Despite the differences between the various approaches proposed (Belcher et al., 2003; Poggi et al., 2004; Okamoto & Nezu, 2009), there is consensus in separating the flow in the canopy from that outside, and in isolating the thin shear layer region at the canopy tip. In the previous subsection we characterised the inner and the outer flow separately; here, instead, we aim at elucidating how they interact with each other. As extensively discussed in literature, the flows over plant canopies (Finnigan, 2000) along with those above permeable (Breugem & Boersma, 2005) and elastic (Rosti & Brandt, 2017) walls are often characterised by the presence of large scale vortical structures elongated in the spanwise direction (rollers), generated by a Kelvin-Helmholtz like instability. We provide evidence of the presence of these structures in our setup in appendix A. These structures are likely to dominate the momentum exchange between the inner and the outer flow, but they are not the only process active close to the shear layer: also high/low speed streaks (Monti et al., 2023) and quasi-streamwise vortices might affect the interaction. The flow dynamics close to the canopy tip is in facts a highly non-trivial phenomenon, which can not be exhaustively described only accounting for the presence of the rollers. Furthermore, as noted by Nicholas et al. (2023), drag model-based simulations provide reasonable results in capturing the flow behaviour near the canopy, but they fall short in providing a detailed characterisation of the flow within the canopy layer and at the interface with the outer flow. A better understanding is therefore crucial for the development of more accurate models.
Thus, in figure 11 we observe the sweep ($u^{\prime}>0,v^{\prime}<0$) and ejection ($u^{\prime}<0,v^{\prime}>0$) events taking place at three locations distinctive of the three regions mentioned above [^1] The continuous evolution of sweeps and ejections along the wall-normal direction can be found in Poggi et al. (2004), for a rigid canopy.: at $y=H/2$, at the canopy tip, corresponding to the outer inflection point $y_{out}$, and at the virtual origin $y_{vo}$. We consider the rigid canopy case (panels $a,b,c$) as well as a flexible one ($Ca=100$, panels $d,e,f$). The most intense sweeps and ejections occur at the canopy tip (panels $b,e$), while they attenuate and become more coherent moving in the outer flow (panels $a,d$). Consistently with the observations of Gao et al. (1989), sweep and ejections are of about equal strength at approximatively twice the canopy height. At the virtual origin (panels $c,f$), instead, only the most intense sweep events are able to penetrate and they occupy small, well defined regions; ejections there are less frequent. Overall, the turbulent state is more strongly driven in the rigid canopy case, where it is characterised by more coherent events compared to the flexible one. Sweep events reach the virtual origin of the flexible canopy flow more frequently, but with diminished intensity.
@@
In order to provide a more quantitative description of the momentum transfer between the inner and the outer flows, we compute the joint probability density function (J-PDF) of the streamwise and wall-normal velocity fluctuations at the canopy tip for different values of $Ca$. As visible in figure 12, the peaks always lay in the second quadrant ($u^{\prime}<0,v^{\prime}>0$) denoting the dominance of ejections over sweeps, as well documented in literature (Gao et al., 1989; Finnigan, 2000). The downward motion of high-speed fluid is mitigated by the presence of the canopy, which instead opposes little resistance to the uplift of low-speed fluid from its interior. Such observation is consistent with the experiment of Chowdhuri et al. (2022), who noted that long-lasting ejection events are more relevant than short-lived sweep events close to the canopy tip. Furthermore, the shape of the J-PDFs closely resembles that reported by Manes et al. (2011) in the case of a highly-permeable porous medium. Notwithstanding this similarity, our J-PDFs differ from one another due to the effect of canopy flexibility. In the most rigid cases the J-PDFs elongate in the fourth quadrant ($u^{\prime}>0,v^{\prime}<0$), remarking the occurrence of rare and intense sweep events which are attenuated increasing the flexibility. In the most flexible cases, in facts, the filaments are deflected by the mean flow and shield the inner region of the canopy from downward interactions. Furthermore, at $Ca=0$ and $Ca=1$, the J-PDFs exhibit a secondary peak in the third quadrant ($u^{\prime}<0,v^{\prime}<0$) associated with the deflection of the streamlines around the filament tip. As seen in Monti et al. (2020) and in figure 16 of Nicholas et al. (2023), immediately after the tip the flow decelerates and plunges.
@@
Our analysis appears to support a scenario where turbulence inside the canopy is sustained by intense events induced by the outer flow. Nevertheless, while the sweeps are hindered by the presence of the filaments, the canopy opposes little resistance to ejections. The filaments, in facts, regulate the vertical exchange of streamwise momentum allowing for the transpiration of the canopy and shielding the inner flow from all but the most intense downward interactions (likely associated to the secondary instability of the rollers, Nepf, 2012a). Such intense sweeps induce a significant local deflection of the filaments and a large-amplitude flapping motion. Turbulence inside the canopy is therefore driven by frequent ejections and rare intense sweep events (Nepf, 2012a), along with the flapping motion of the filaments in the flexible cases.